OpenStudy (anonymous):

Let I be a bounded closed interval, and let f : I -> I be continuous. Show that there exists x in I such that f(x) = x. Are the hypotheses about I both necessary?

OpenStudy (anonymous):

@Hero help?

OpenStudy (anonymous):

@Zarkon please help?

OpenStudy (anonymous):

I don't know where to start

terenzreignz (terenzreignz):

Mind you, Real analysis isn't my best field... let's put together stuff that we know... One, it's continuous, and on a closed and bounded interval... that means f has an absolute max and absolute min on that interval, right?

OpenStudy (anonymous):

thanks! and yes.

terenzreignz (terenzreignz):

Actually, there may be a slight problem...

OpenStudy (anonymous):

I did some research and I'm wondering if this would help? http://www.physicsforums.com/showthread.php?t=575366

terenzreignz (terenzreignz):

hang on :D

OpenStudy (john_es):

May be there is a possibility to find a proof using the intermediate value theorem, without sequences.

terenzreignz (terenzreignz):

My sentiments exactly :D Just finding out how...

terenzreignz (terenzreignz):

Okay, I got it now :)

OpenStudy (anonymous):

ok i'm going to try that

terenzreignz (terenzreignz):

Oh... sure... let me know if you get stuck :D

OpenStudy (anonymous):

i actually used the fact that since f is continuous, there would exist a sequence \((x_n)\) in I with \(\lim x_n=c\) and the sequence \((f(x_n))\) converges to \(f(c)\)

terenzreignz (terenzreignz):

That's messy don't you think? :)

OpenStudy (anonymous):

that was a property given in the text though

terenzreignz (terenzreignz):

It was specifically mentioned that you are to use the Sequential Criterion?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

i had the idea after reading that forum.

terenzreignz (terenzreignz):

Then hear me out :)

OpenStudy (anonymous):

Also, my professor hadn't gone over IVT yet

terenzreignz (terenzreignz):

Oh... that is a problem...

terenzreignz (terenzreignz):

But he's already gone through the sequential criterion?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

Interesting... let's rethink this...

OpenStudy (anonymous):

I have \(x_1=a_0\) and \(x_{n+1}=f(x_n)\) and then I take the limit on both sides of the second equation

OpenStudy (anonymous):

I set \(lim(x_n)=c\) because it's continuous, then I got \(c=f(c)\) but I'm not sure if it would be so easy

terenzreignz (terenzreignz):

Frustrating... IVT gets this done in nothing flat...

OpenStudy (anonymous):

how do i do it with IVT though?

OpenStudy (anonymous):

I'll just explain to him what I would do

terenzreignz (terenzreignz):

It's actually really simple... let I = [a,b] (the closed and bounded interval)

terenzreignz (terenzreignz):

Then we have f(a) and f(b)

terenzreignz (terenzreignz):

Inevitably, one of these would be greater than the other, so let's assume f(b) > f(a)

OpenStudy (anonymous):

ok, i understand that part.

terenzreignz (terenzreignz):

Wait.... hold that thought... wrong assumption... rewind... start over (sorry, not myself tonight :3)

OpenStudy (anonymous):

it's ok

terenzreignz (terenzreignz):

Since I is closed and bounded, f attains an absolute minimum for some b in I and an absolute minimum for some a in I, aye? (LOL)

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

I meant an absolute maximum for some b.

OpenStudy (anonymous):

lol

terenzreignz (terenzreignz):

Wait something's wrong again... gah... never mind, false alarm :/

OpenStudy (anonymous):

he just sent out an email saying we're allowed to use IVT but since we haven't gone over it, I am not quite familiar with it to use it

terenzreignz (terenzreignz):

Okay, first, let me introduce it to you :)

terenzreignz (terenzreignz):

In its simplest form we have... suppose we have a continuous function f on a closed and bounded interval [a,b]

terenzreignz (terenzreignz):

Where f(a)f(b) < 0

terenzreignz (terenzreignz):

Catch me so far?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

That is to say, the product of f(a) and f(b) is negative

terenzreignz (terenzreignz):

then there exists a c between a and b such that f(c) = 0

OpenStudy (anonymous):

then there would be a c in which f(c) =0?

OpenStudy (anonymous):

oh wait you already explained it lol

terenzreignz (terenzreignz):

Yup... can you visualize that? :D

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

The meaning of f(a)f(b) < 0 is that f(a) and f(b) must have different signs.

terenzreignz (terenzreignz):

Oh okay, great :D

OpenStudy (anonymous):

it's basically the same application of IVT in calculus, right?

terenzreignz (terenzreignz):

Yup. So it turns out I was right the first time.... I just panicked. :3

OpenStudy (anonymous):

ok, hit me! lol

terenzreignz (terenzreignz):

So, we let I = [a,b] our interval.

terenzreignz (terenzreignz):

And our function is f from I to I.

terenzreignz (terenzreignz):

Well, in that case, first we assume f(a) is not equal to a and f(b) is not equal to b. Of course, if this weren't true, then we'd already be done.

terenzreignz (terenzreignz):

So far so good?

OpenStudy (anonymous):

so we're going to prove by contradiction?

terenzreignz (terenzreignz):

Not exactly... we assume first that f(a) \(\ne\) a and f(b) \(\ne\) b

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Now, since the function's codomain is also I, then a is the lowest value that f could possibly take on that interval.

OpenStudy (anonymous):

ok i'm with you

terenzreignz (terenzreignz):

So since f(a) is not a, it follows that f(a) > a

terenzreignz (terenzreignz):

Similarly, b is the highest value f could possibly take on the interval I, so since f(b) is not b, it follows that g(b) < b

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Good... now define a new function g(x) = f(x) - x

OpenStudy (anonymous):

by g(b) up there, you did mean f(b) right?

terenzreignz (terenzreignz):

whoops... sorry, got a little too excited :) Yeah, f(b)

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

So, we now consider g(a) and g(b) Since f(a) > a, then it follows that f(a) - a = g(a) > 0

terenzreignz (terenzreignz):

And since f(b) < b, then it follows that f(b) - b = g(b) < 0

OpenStudy (anonymous):

ow do we know g(a) would be >0?

terenzreignz (terenzreignz):

Because we know that f(a) > a

OpenStudy (anonymous):

ohh ok thanks i was stupid

terenzreignz (terenzreignz):

from assuming that f(a) is not equal to a, and knowing that a is the lowest value that f could possibly take on the interval I.

terenzreignz (terenzreignz):

So, we now have g(a) > 0 g(b) < 0

terenzreignz (terenzreignz):

And mind you, g is continuous, being the difference of two continuous functions.

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

That now means g(a)g(b) < 0

terenzreignz (terenzreignz):

Implying there is a value x somewhere between a and b where g(x) = 0

terenzreignz (terenzreignz):

Or f(x) - x = 0

terenzreignz (terenzreignz):

And the rest is history :)

OpenStudy (anonymous):

YAY you're awesome.

OpenStudy (anonymous):

How do I know if these hypotheses were necessary, though?

terenzreignz (terenzreignz):

We needed the interval I to be bounded... otherwise it could be that I is the entire set of Real numbers

terenzreignz (terenzreignz):

In which case, an easy counterexample would be f(x) = x+1 Which has no value of x such that f(x) = x And yet f is a continuous function from I to I.

terenzreignz (terenzreignz):

So we needed it to be a bounded interval... we needed it to be continuous on the CLOSED interval [a,b] because we needed f(a) and f(b)

OpenStudy (anonymous):

ok i got it! thanks so much!

terenzreignz (terenzreignz):

sure :)